1) Please answer full question (a-c) showing work. Ensure answer is correct for thumbs up.

a) If the FFT of an 1000-Hz square wave was measured to be approximately -1.8 dBV. What is the corresponding rms value?

b) The third harmonic of the square wave was measured to be approximately -11.34 dBV. What is the corresponding rms value?

c) The output voltage across an 8-Ω speaker was measured as 1-kHz sinusoidal signal with peak

amplitude 7 Vp. Calculate the value of the power delivered to the speaker in watts? Express your result in dBW.
Answer #1

(a) If the FFT of an 1000-Hz square wave was measured to be approximately -1.8 dBV. What is the corresponding rms value?

ANSWER:

given VdBV= -1.8

=> 20 log10 Vrms = -1.8

=> log10 Vrms = -1.8/20 = -0.09

=> Vrms = 10-0.09=0.81 VOLT

b) The third harmonic of the square wave was measured to be approximately -11.34 dBV. What is the corresponding rms value?

ANSWER:

Given

VdBV= -11.34

=> 20 log10 Vrms = -11.34

=> log10 Vrms = -11.34/20 = -0.567

=> Vrms = 10-0.567= 0.27 VOLT

c) The output voltage across an 8-Ω speaker was measured as 1-kHz sinusoidal signal with peak amplitude 7 Vp. Calculate the value of the power delivered to the speaker in watts? Express your result in dBW.

ANSWER: